OpenStudy (errinss):

Solve the equation. x2 = 96

pooja195 (pooja195):

Think back to the one we did previously

OpenStudy (cas_fangirl_14):

\[96\div2=?\]

pooja195 (pooja195):

^nope \[\huge~\rm~\bf x^2=96\]

OpenStudy (errinss):

we would square root x and 96 right?

pooja195 (pooja195):

Yep :)

OpenStudy (cas_fangirl_14):

I didn't know it was ^2

OpenStudy (errinss):

x=9.7979589711327123927891362988236

satellite73 (satellite73):

yo no

pooja195 (pooja195):

break down 96

OpenStudy (errinss):

a:9,-9 b:9 c d are the following pictures

OpenStudy (errinss):

those are my choices for this problem

satellite73 (satellite73):

\[x^2=96 \\ x=\pm\sqrt{96}\] is a star

pooja195 (pooja195):

^as I said, you would break down 96.

OpenStudy (errinss):

lol thanks

OpenStudy (cheez):

No prob. Glad I could help again.

OpenStudy (errinss):

again...hehe thx X'D